# A Unifying Framework for Parallelizing Sequential Models with Linear Dynamical Systems

### v1

@@ 2.3 Picard iterations
A seemingly different approach to using fixed-point iterations are Picard iterations, which were used by Shih et al. (2023) to parallelize sampling in diffusion models. Picard iterations are often used in the context of evaluating ODEs, where
$\dot{x}=g(x,t).$ (7)
After Euler discretization with step size $\Delta$, the continuous time eq.˜7 becomes the discrete-time recursion,
@@
Shih et al. (2023) proves by induction that for any dynamical system given by eq.˜8, the fixed-point iterations given by eq.˜9 will converge to the true trajectory in at most $T$ iterations. Similarly, Gonzalez et al. (2024) proved that for any dynamical system given by eq.˜1, both the Newton fixed-point iterations given by eq.˜5 and the quasi-Newton fixed-point iterations given by eq.˜6 will converge to the true trajectory in at most $T$ iterations. The similarity of these results and techniques begs the question as to how Picard and Newton iterations relate to each other. Our first result shows that Picard iterations are in fact a type of quasi-Newton iteration, where we approximate the Jacobian of the dynamics function by the identity matrix.
@@
The Picard iteration operator $\mathcal{A}_{P}$ given by eq.˜9 is a special case of an LDS, eq.˜4, where the transition matrix is the identity,
$A_{t+1}=I_{D}.$
@@
Define $f_{t+1}(x_{t}):=x_{t}+g(x_{t},t)\cdot\Delta$. Then, from eq.˜9 it follows that
@@
This is exactly of the form of the generic linear recursion shown in eq.˜4, with $A_{t+1}=I_{D}$. ∎
An important consequence of Proposition˜1 is that like Newton iterations and quasi-Newton iterations, Picard iterations can also be cast as an LDS. In Newton iterations, the full Jacobian $\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}$ is used in LDS; in quasi-Newton iterations, the diagonal approximation $\mathrm{diag}[\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}]$ is used; and in Picard iterations, the identity $I_{D}$ is used. The Picard iteration is more compute and memory efficient than even quasi-Newton, but is also generally a less faithful approximation and takes more iterations to converge, unless the Jacobian is well-approximated by the identity.
@@ 2.4 Jacobi iterations
Yet another seemingly different fixed-point method are Jacobi iterations (Ortega & Rheinboldt, 1970), which were used by Song et al. (2021a) to accelerate computation in a variety of settings in machine learning, such as feedforward networks with skip connections. Jacobi iterations are also a zeroth-order fixed-point method, and are commonly used to solve systems of multivariate nonlinear equations of the form,
@@
Song et al. (2021a) considers in particular the problem of solving recurrence relations of the form $x_{t+1}=f_{t+1}(\mathbf{x}_{1:t})$}, and proves that, for such a system, Jacobi iterations converge in at most $T$ iterations. This result is directly analogous to similar such results and proofs in Gonzalez et al. (2024) for Newton and quasi-Newton iterations and Shih et al. (2023) for Picard iterations. However, as Song et al. (2021a) notes, for a Markovian system (the setting we consider in this paper), it takes $T$ Jacobi iterations for information from $x_{0}$ to propagate to $x_{T}$. Thus, Jacobi iterations are not generally suitable for Markovian processes like state space models, RNNs, or sampling from a diffusion model. In fact, in the context of iteratively applying LDSs to parallelize Markovian state space models, we prove that Jacobi iterations are a type of degenerate quasi-Newton iterations, where we “approximate” the Jacobian of the dynamics function by zero.
@@
When applied to a Markovian state space model as in eq.˜1, the Jacobi iteration operator $\mathcal{A}_{J}$ specified by eq.˜10 is a special case of the common form, eq.˜4, where,
$A_{t+1}=0.$
@@
2.5 Summary
We have shown how important parallel fixed-point iterations—Newton, quasi-Newton, Picard, and Jacobi iterations—can all be cast as LDSs when deployed for evaluating nonlinear recursions, as summarized in Table˜1. The regimes where these different methods excel are therefore dictated by the form of the Jacobians of their dynamics functions: if each $f_{t+1}$ is close to an identity update (as is the case in sampling from a diffusion model with small discretization parameter), then Picard will excel; if the dynamics are nearly uncoupled across state dimensions, then quasi-Newton using a diagonal approximation will excel; and if the dynamics have multiple dependencies across coordinates and the dimension $D$ is not too large, then Newton iterations will excel. On a Markovian state space model, Jacobi iterations require $T$ iterations for information to propagate from the initial condition to the end of the sequence, the same as sequential evaluation. However, we include Jacobi iterations for completeness, noting that they can be useful in non-Markovian systems, such as architectures with skip connections.
An important corollary is that because all of these fixed-point iterations can be cast as LDSs, they are all guaranteed to converge in all problem settings in at most $T$ iterations (Gonzalez et al., 2024, Proposition 1). However, as we noted above, the precise convergence rates of the different fixed-point methods will be problem dependent. In the next section, we develop these observations further by discussing how a wide variety of important tasks in machine learning can be understood as the evaluation of Markovian state space models. Furthermore, we show how the unifying framework proposed in this section helps us understand which fixed-point iterations will excel in which problem settings.
3 Case studies on when to use each kind of fixed-point iteration
Many important tasks in machine learning can be understood as the evaluation of a sequential process. For example, evaluating deep neural networks, recurrent neural networks, state space models, or transformers (sequential evaluation of the blocks over depth); sampling from a diffusion model, or in general any Markov chain or process such as Markov chain Monte Carlo; maintaining a world model over the course of many sequential updates to the state; and many more problems fall under this heading. In this section, we consider three empirical case studies that illustrate how the unifying framework above provides heuristic guidance about which fixed-point schemes will excel in which settings. This concordance is based on the structure of the Jacobian of $f_{t+1}$ and the relative computational cost of different fixed-point methods.
Our heuristic guidance, in a nutshell, is:
@@ 3 Case studies on when to use each kind of fixed-point iteration
Simpler approximate Jacobians, like the identity matrix Picard iterations, are less computationally expensive, meaning that each fixed-point iteration is more efficient. So, if the lower-order fixed-point method still converges in a small number of fixed-point iterations, it achieves the sequential roll-out $\mathbf{x}^{*}$ in faster wall-clock times on GPUs than higher-order fixed point methods. However, if the higher-order fixed-point method (e.g. Newton or quasi-Newton) converges in far fewer iterations than the lower-order fixed-point method, then the increased computation of the higher-order fixed-point method is worthwhile. Intuitively, the number of iterations needed for a fixed-point method to converge should be related to how faithfully $A_{t}$ approximate the linearized dynamics $\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}$. We support this intuition with three empirical case studies to highlight where Newton, quasi-Newton, and Picard iterations excel. Further experimental details are provided in Appendix˜C. Our code is available at https://github.com/lindermanlab/parallelizing_with_lds
@@ 3.1 Case study #1: Solving the group word problem with Newton iterations
Newton iterations should outperform quasi-Newton and Picard iterations in settings where the Jacobian of the recursion, $f_{t+1}$, is not well approximated by its diagonal or by the identity matrix. One example of such a recursion is the group word problem, which has been used to theoretically and empirically assess the limits of sequential modeling architectures for state-tracking tasks (Kim & Schuster, 2023; Liu et al., 2023; Merrill et al., 2024; Grazzi et al., 2025; Schöne et al., 2025). In the sequence-modeling community, the term “group word problem“ is defined as follows.
@@
Merrill et al. (2024) emphasizes that nonlinear RNNs in both theory and practice are able to learn the group word problem in arbitrary groups to high accuracy with only a single layer, whereas compositions of popular linear RNNs linked by nonlinearities, such as S4 (Gu et al., 2022) and Mamba [^2] Mamba allows input-dependent dynamics matrices but they must be diagonal, which prevents a single Mamba layer from implementing the particular LDS in Proposition 3, which uses permutation matrices. (Gu & Dao, 2024), require a number of layers that grows with $T$. Merrill et al. (2024) emphasizes that recurrent architectures with nonlinear transitions are well-suited for solving the group word problem, because in theory and practice, such architectures can learn the group word problem to high accuracy with a single layer. Other literature has explored the value of matrix-valued states (Beck et al., 2024; Grazzi et al., 2025). However, in Proposition˜3 below, we show that neither nonlinearity nor matrix-valued states are needed to understand or solve the group word problem. Instead, the problem can be formulated as an LDS with vector-valued states and input-dependent transition matrices.
@@
By Cayley’s theorem, any finite group $G$ can be embedded in a symmetric group $S_{D}$, for some $D\leq|G|$. Therefore, by choosing the initial state $x_{0}\in\mathbb{R}^{D}$ to have $D$ distinct entries (a “vocabulary” of size $D$), we can use the tabular representation of permutations (Artin, 2011, eq. 1.5.2) to represent an element of $S_{D}$ as $x_{t}$ (by a permutation of the elements of $x_{0}$). We can also choose $A_{t+1}\in\mathbb{R}^{D\times D}$ to be the permutation matrix corresponding to the embedding of $g_{t+1}$ in $S_{D}$, since any element of $S_{D}$ can be represented as a $D\times D$ permutation matrix (e.g., see Figure˜1B). Consequently $x_{t}=A_{t}A_{t-1}\ldots A_{2}A_{1}x_{0}$ is an embedding of an element of $G$ in $S_{D}$ in the tabular representation. In fact, $x_{t}\in\mathbb{R}^{D}$ represents the running product $g_{1}g_{2}\ldots g_{t-1}g_{t}$, which is precisely the goal of the group word problem. ∎
@@
Figure 1: A single Newton iteration solves the $S_{5}$ group word problem, whereas the number of iterations required for the quasi-Newton and Picard methods increases with problem size. We consider the task of evaluating the product of $S_{5}$ group elements. A: The group word problem can be expressed as an LDS with input-dependent state-transition matrices. B: An example input-dependent transition matrix $A_{t}$ for permutation (1 5 2 4 3)(1\ 5\ 2\ 4\ 3), in cycle notation. C: For each fixed-point method and a range of sequence lengths, $T$, we compute the median (over ten random seeds) number of fixed-point iterations to converge (top) and the median wall-clock time (bottom). While a single Newton iteration is sufficient to solve the $S_{5}$ problem, the number of iterations required for the Picard and quasi-Newton methods increases with the sequence length.
@@
Though we have cast the group word problem as a time-varying LDS with $f_{t+1}(x_{t})=A_{t+1}x_{t}$, note that we can still evaluate this recursion with any of the fixed-point methods described above. Since the dynamics are linear, the Newton iteration corresponds to evaluating the LDS with a parallel scan, and it converges in one iteration. While quasi-Newton or Picard methods would require more iterations to converge, they could still be more efficient in wall-clock time, since they use less memory and compute per iteration. However, since the input-dependent state transition matrices of the group word problem have mostly non-diagonal entries, as they are permutation matrices, we expect the diagonal and identity transition matrices in the quasi-Newton and Picard iterations, respectively, to be poor approximations of the true dynamics, and therefore these methods should take a large number of fixed-point iterations to converge.
@@
We support this hypothesis with a simple experiment simulating the $S_{5}$ word problem, a standard problem in the sequence modeling literature (Merrill et al., 2024; Grazzi et al., 2025). In this setting, Figure˜1 shows that quasi-Newton and Picard iterations require nearly $T$ iterations to converge. On the other hand, we see that Newton’s method solves the $S_{5}$ word problem with just one fixed-point iteration, as expected since the true dynamics are linear. The speed-up is also apparent in the wall-clock time comparison, where we see that Newton is faster than quasi-Newton, Picard, and sequential evaluation, regardless of $T$.
3.2 Case Study #2: Parallelizing RNNs with quasi-Newton iterations

### v2

@@ 2.3 Picard iterations
A seemingly different approach to using fixed-point iterations is Picard iterations, which were used by Shih et al. (2023) to parallelize sampling in diffusion models. Picard iterations are often used in the context of evaluating ODEs, where
$\dot{x}=g(x,t).$
After Euler discretization with step size $\Delta$, we obtain the discrete-time recursion,
@@
Shih et al. (2023) proves by induction that for any dynamical system given by eq.˜7, the fixed-point iterations given by eq.˜8 will converge to the true trajectory in at most $T$ iterations. Similarly, Gonzalez et al. (2024) proved that for any dynamical system given by eq.˜1, both the Newton fixed-point iterations given by eq.˜5 and the quasi-Newton fixed-point iterations given by eq.˜6 will converge to the true trajectory in at most $T$ iterations. The similarity of these results and techniques begs the question as to how Picard and Newton iterations relate to each other. Our first result shows that Picard iterations are in fact a type of quasi-Newton iteration, where we approximate the Jacobian of the dynamics function by the identity matrix.
@@
The Picard iteration operator $\mathcal{A}_{P}$ given by eq.˜8 is a special case of an LDS, eq.˜4, where the transition matrix is the identity,
$\tilde{A}_{t+1}=I_{D}.$
@@
Define $f_{t+1}(x_{t}):=x_{t}+g(x_{t},t)\cdot\Delta$. Then, from eq.˜8 it follows that
@@
This is exactly of the form of the generic linear recursion shown in eq.˜4, with $\tilde{A}_{t+1}=I_{D}$. ∎
An important consequence of Proposition˜1 is that like Newton iterations and quasi-Newton iterations, Picard iterations can also be cast as an LDS. In Newton iterations, the full Jacobian $\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}$ is used in LDS; in quasi-Newton iterations, the diagonal approximation $\mathrm{diag}[\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}]$ is used; and in Picard iterations, the identity $I_{D}$ is used. The Picard iteration is more compute and memory efficient than even quasi-Newton, requiring only vector additions via the prefix sum algorithm. However, it is generally a less faithful approximation and takes more iterations to converge, unless the Jacobian is well-approximated by the identity.
@@ 2.4 Jacobi iterations
Yet another seemingly different class of fixed-point methods are Jacobi iterations (Ortega and Rheinboldt, 1970), which were used by Song et al. (2021a) to accelerate computation in a variety of settings in machine learning, such as feedforward networks with skip connections. Jacobi iterations are also a zeroth-order fixed-point method, and are commonly used to solve systems of multivariate nonlinear equations of the form,
@@
Song et al. (2021a) considers in particular the problem of solving recurrence relations of the form $x_{t+1}=f_{t+1}(\mathbf{x}_{1:t})$}, and proves that, for such a system, Jacobi iterations converge in at most $T$ iterations. This result is directly analogous to similar such results and proofs in Gonzalez et al. (2024) and Tang et al. (2024) for Newton and quasi-Newton iterations and Shih et al. (2023) for Picard iterations. In fact, in the context of iteratively applying LDSs to parallelize Markovian state space models, we prove that Jacobi iterations are a type of degenerate quasi-Newton iterations, where we “approximate” the Jacobian of the dynamics function by zero.
@@
When applied to a Markovian state space model as in eq.˜1, the Jacobi iteration operator $\mathcal{A}_{J}$ specified by eq.˜9 is a special case of the common form, eq.˜4, where,
$\tilde{A}_{t+1}=0.$
@@
We have shown that evaluating nonlinear recursions via Newton, quasi-Newton, Picard, and Jacobi fixed-point iterations reduces to iteratively solving linear dynamical systems. An important corollary is that these methods are guaranteed to converge in all problem settings in at most $T$ iterations (Gonzalez et al., 2024, Prop. 1). However, the precise convergence rates of the different fixed-point methods will depend on the accuracy of the Jacobian approximation, as we show next.
3 Analysis of convergence rates
In this section, we analyze the convergence properties of the fixed-point methods above. We show that the convergence rate of these fixed-point methods depends on two factors: (i) how well the transition matrix $\tilde{A}_{t+1}$ approximates the true dynamics Jacobian $A_{t+1}:=\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}$; and (ii) the stability of the LDS with transition matrices $\tilde{A}_{1:T}$.
3.1 Derivation of fixed-point iterations from high-dimensional root-finding.
We first review the derivation of the Newton fixed-point iterations in eq.˜5 from the system of equations given by eq.˜3, highlighting how the other fixed-point iterations naturally arise from different approximations of $\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}.$ This background will be useful for proving convergence rates in Section˜3.2.
Returning to the system of nonlinear equations in eq.˜3, we define the residual function $\mathbf{r}:\mathbb{R}^{TD}\mapsto\mathbb{R}^{TD}$,
$\mathbf{r}(\mathbf{x}_{1:T})=[x_{1}-f_{1}(x_{0}),\,x_{2}-f_{2}(x_{1}),\,x_{3}-f_{3}(x_{2}),\ldots,x_{T}-f_{T}(x_{T-1})].$ (10)
We define $\mathbf{J}(\mathbf{x}_{1:T})\in\mathbb{R}^{TD\times TD}$ to be the Jacobian of the residual $\mathbf{r}(\mathbf{x}_{1:T})$,
$\mathbf{J}(\mathbf{x}_{1:T}):=\dfrac{\partial\mathbf{r}}{\partial\mathbf{x_{1:T}}}(\mathbf{x}_{1:T})=\begin{pmatrix}I_{D}&0&0&\ldots&0&0\\ -\frac{\partial f_{2}}{\partial x}(x_{1})&I_{D}&0&\ldots&0&0\\ 0&-\frac{\partial f_{3}}{\partial x}(x_{2})&I_{D}&\ldots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\ldots&I_{D}&0\\ 0&0&0&\ldots&-\frac{\partial f_{T}}{\partial x}(x_{T-1})&I_{D}\\ \end{pmatrix}$. (11)
Therefore, in this notation, every Newton’s method update takes the form
$\mathbf{x}_{1:T}^{(i+1)}=\mathcal{A}_{N}(\mathbf{x}_{1:T}^{(i)}):=\mathbf{x}_{1:T}^{(i)}-\mathbf{J}(\mathbf{x}_{1:T}^{(i)})^{-1}\mathbf{r}(\mathbf{x}_{1:T}^{(i)}).$
Using matrix manipulations, it can be shown (Danieli et al., 2023; Lim et al., 2024; Gonzalez et al., 2024) that each Newton fixed-point iteration is given by the linear recursion in eq.˜5.
Analogously, for the other fixed point iterations considered in this paper (cf. Table˜1), which use transition matrices $\tilde{A}_{t+1}$ to approximate $\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}$, we can define a block-bidiagonal $\widetilde{\mathbf{J}}\in\mathbb{R}^{TD\times TD}$ simply by substituting $\tilde{A}_{t+1}(x_{t})$ for $\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}$ in eq.˜11. The corresponding fixed point iteration $\mathcal{A}$ takes the form
$\mathcal{A}(\mathbf{x}_{1:T}^{(i)}):=\mathbf{x}_{1:T}^{(i)}-\widetilde{\mathbf{J}}(\mathbf{x}_{1:T}^{(i)})^{-1}\mathbf{r}(\mathbf{x}_{1:T}^{(i)}).$ (12)
Different fixed-point methods $\mathcal{A}$ give rise to different matrices $\widetilde{\mathbf{J}}$, which impacts their convergence rates.
3.2 Convergence rates of fixed-point iterations
We derive convergence rates for all fixed-point operators discussed in this paper. The analysis follows standard techniques in the literature (cf. Kelley (1995, Thm 5.4.1)), which Lu et al. (2025, Prop 4.) applied to parallelizing sequential models with Picard iterations. Our unifying framework allows us to see that their analysis generalizes straightforwardly.
Let $\mathbf{e}^{(i)}_{1:T}:=\mathbf{x}_{1:T}^{(i)}-\mathbf{x}_{1:T}^{\star}\in\mathbb{R}^{TD}$ denote the error of a trajectory $\mathbf{x}_{1:T}^{(i)}$. Note that the error depends on the current trajectory and the dynamics functions, but we suppress these dependencies to keep notation lightweight. For any of the fixed-point methods discussed in this paper, we can bound the convergence rate of this error.
Proposition 3 (c.f. Proposition 4 of Lu et al. (2025)).
Consider a fixed-point solver with updates given by eq.˜12 for some matrix $\widetilde{\mathbf{J}}(\mathbf{x}_{1:T}^{(i)})$ with form specified by eq.˜19. Let $L$ be the maximum of the Lipschitz constants of $\nicefrac{{\partial f_{t}}}{{\partial x_{t-1}}}$. Then $\|\mathbf{e}^{(i+1)}_{1:T}\|_{2}$ satisfies
$\|\mathbf{e}^{(i+1)}_{1:T}\|_{2}\leq\left\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T}^{(i)})^{-1}\right\|_{2}\cdot\left(\left\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T}^{(i)})-\mathbf{J}(\mathbf{x}_{1:T}^{(i)})\right\|_{2}\|\mathbf{e}^{(i)}_{1:T}\|_{2}+\frac{L}{2}(\|\mathbf{e}^{(i)}_{1:T}\|_{2}^{2})\right),$ (13)
where $\|\cdot\|_{2}$ denotes the spectral norm of a matrix and the $\ell_{2}$ norm of a vector.
Proof.
Starting from eq.˜12, we subtract $\mathbf{x}_{1:T}^{\star}$ from both sides to obtain
$\mathbf{e}^{(i+1)}_{1:T}=\mathbf{e}^{(i)}_{1:T}-\widetilde{\mathbf{J}}(\mathbf{x}_{1:T}^{(i)})^{-1}\mathbf{r}(\mathbf{x}_{1:T}^{(i)}).$
Next, we Taylor expand $\mathbf{r}(\cdot)$ around $\mathbf{x}_{1:T}^{(i)}$ to obtain
$\mathbf{r}(\mathbf{x}_{1:T}^{\star})=\mathbf{r}(\mathbf{x}_{1:T}^{(i)})-\mathbf{J}(\mathbf{x}_{1:T}^{(i)})\mathbf{e}^{(i)}_{1:T}+\mathbf{R}(\mathbf{e}^{(i)}_{1:T})$,
where $\mathbf{R}(\mathbf{e}^{(i)}_{1:T})$ is the second-order remainder function and has norm bounded by $\nicefrac{{\|\mathbf{e}^{(i)}_{1:T}\|_{2}^{2}}}{{2}}$ times the Lipschitz constant of $\mathbf{J}(\mathbf{x}_{1:T}^{(i)})$, which Theorem 3 of Gonzalez et al. (2025) shows is bounded by $L$. Since $\mathbf{r}(\mathbf{x}_{1:T}^{\star})=\mathbf{0}$, it follows that
$\mathbf{e}^{(i+1)}_{1:T}=\widetilde{\mathbf{J}}(\mathbf{x}_{1:T}^{(i)})^{-1}\bigg(\left(\widetilde{\mathbf{J}}(\mathbf{x}_{1:T}^{(i)})-\mathbf{J}(\mathbf{x}_{1:T}^{(i)})\right)\mathbf{e}^{(i)}_{1:T}+\mathbf{R}(\mathbf{e}^{(i)}_{1:T})\bigg)$. (14)
The result follows by taking norms on both sides and using the triangle inequality. ∎
3.3 Intuitions about rates of convergence
Proposition˜3 shows that, to first order, the error is determined by the discrepancy between the chosen linear operator $\widetilde{\mathbf{J}}$ and the true Jacobian $\mathbf{J}$ of the residual. Moreover, we see that as $\|\mathbf{e}^{(i)}_{1:T}\|_{2}$ approaches zero, the first-order term, which is linear in $\|\mathbf{e}^{(i)}_{1:T}\|_{2}$, must eventually (under strong enough continuity assumptions) dominate the second-order term, which is quadratic in $\|\mathbf{e}^{(i)}_{1:T}\|_{2}$. Typically, we would say the rate of decrease in $\|\mathbf{e}^{(i)}_{1:T}\|_{2}$ approaches an asymptotic linear rate $\gamma$ given by
$\gamma:=\left\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T}^{\star})^{-1}\right\|_{2}\left\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T}^{\star})-\mathbf{J}(\mathbf{x}_{1:T}^{\star})\right\|_{2}.$ (15)
Discussions of asymptotic linear rate are subtle in our setting, where all fixed-point methods are guaranteed to converge in $T$ iterations (see our discussion in Section˜F.1). Nonetheless, the functional form of $\gamma$ provides useful intuition about the convergence rates of different fixed-point methods.
Let’s study the two factors in the asymptotic linear rate $\gamma$, which must be less than one to be meaningful. Lu et al. (2025) left quantifying these two factors for future work, which we now bring to fruition. The first factor is the spectral norm of the inverse of the approximate Jacobian. Because $\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})$ is a block bidiagonal matrix (see eq.˜11 and the definition that follows), its inverse has a block lower triangular structure of the form
$\widetilde{\mathbf{J}}(\mathbf{x}_{1:4})^{-1}=\begin{pmatrix}I_{D}&0&0&0\\ \tilde{A}_{2}&I_{D}&0&0\\ \tilde{A}_{3}\tilde{A}_{2}&\tilde{A}_{3}&I_{D}&0\\ \tilde{A}_{4}\tilde{A}_{3}\tilde{A}_{2}&\tilde{A}_{4}\tilde{A}_{3}&\tilde{A}_{4}&I_{D}\end{pmatrix}$, (16)
shown above for $T=4$.
From eq.˜16, we see that the blocks of $\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})^{-1}$ are products of the transition matrices $\tilde{A}_{t+1}$ from the chosen fixed-point method. In particular, if the chosen fixed point method results in an unstable LDS with $\|\tilde{A}_{t+1}\|_{2}>1$ for many time points, the spectral norm of $\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})^{-1}$ can be much larger than one. Intuitively, fixed-point methods resulting in unstable LDSs will have slower rates of convergence since instability leads to numerical divergence. We elaborate on this point in Section˜C.3.2.
The second factor is the spectral norm of the difference between the approximate and true residual Jacobian. The following lemma shows that we can control this factor in terms of the spectral norms of the differences between the approximate and true dynamics Jacobians. Intuitively, the rate of convergence will be faster if the transition matrices $\tilde{A}_{t+1}$ are closer to the true dynamics Jacobian $A_{t+1}:=\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}$ in spectral norm.
Lemma 1.
If $\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})$ is given by eq.˜19 and $\mathbf{J}(\mathbf{x}_{1:T})$ is given by eq.˜11, then
$\mathrm{diff}(\mathcal{A}):=\left\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})-\mathbf{J}(\mathbf{x}_{1:T})\right\|_{2}=\max_{1\leq t\leq T-1}\left\|\tilde{A}_{t+1}(x_{t})-A_{t+1}(x_{t})\right\|_{2},$
where $A_{t+1}:=\frac{\partial f_{t+1}}{\partial x_{t}}(x_{t})$.
Proof.
Proposition˜3 is an upper bound on the norm of the error of each fixed-point iterate. As an upper bound, this result can not always fully predict the precise trajectory of the norm of the error (see Section˜C.2). Nevertheless, this proposition provides helpful intuitions that are borne out in the case studies below.
4 Case studies: performance of the different fixed-point methods
In this section, we consider three empirical case studies that illustrate how the unifying framework and convergence analysis presented in this paper provides guidance about which fixed-point schemes will excel in which settings. This concordance is based on the structure of the Jacobian of $f_{t+1}$ and the relative computational cost of different fixed-point methods. In a nutshell, we advise:
@@ 4 Case studies: performance of the different fixed-point methods
To elaborate: simpler approximate Jacobians are less computationally expensive, meaning that each fixed-point iteration is more efficient. So, if a lower-order fixed-point method (e.g., Picard or Jacobi) converges in a small number of iterations, it may reach $\mathbf{x}^{\star}$ in less time than higher-order methods. However, if the higher-order fixed-point method (e.g. Newton or quasi-Newton) converges in far fewer iterations than lower-order methods, then the increased computation of the higher-order fixed-point method may be worthwhile. As supported by the theoretical analysis in Section˜3, the number of iterations needed for a fixed-point method to converge is related to the difference in spectral norm between $\tilde{A}_{t}$ and $A_{t}:=\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}$. We support this intuition with the following case studies, with further experimental details provided in Appendix˜E. Our code is available at https://github.com/lindermanlab/parallelizing_with_lds
@@ 4.1 Case study #1: Solving the group word problem with Newton iterations
Newton iterations should outperform quasi-Newton and Picard iterations in settings where the Jacobian of the recursion, $f_{t+1}$, is not well approximated by its diagonal, the identity matrix, or the zero matrix. One example of such a recursion is the group word problem, which has been used to theoretically and empirically assess the limits of sequential modeling architectures for state-tracking tasks (Kim and Schuster, 2023; Liu et al., 2023; Merrill et al., 2024; Grazzi et al., 2025; Schöne et al., 2025). In the sequence-modeling community, the term “group word problem” is defined as follows.
@@
Merrill et al. (2024) emphasizes that nonlinear RNNs in both theory and practice are able to learn the group word problem in arbitrary groups to high accuracy with only a single layer, whereas compositions of popular linear RNNs linked by nonlinearities, such as S4 (Gu et al., 2022) and Mamba [^3] Mamba allows input-dependent dynamics matrices but they must be diagonal, which prevents a single Mamba layer from implementing the particular LDS in Proposition 4, which uses permutation matrices. Merrill et al. (2024) also demonstrate that a linear time-varying system with a dense transition matrix can learn the group word problem.  (Gu and Dao, 2024), require a number of layers that grows with $T$. Merrill et al. (2024) emphasizes that recurrent architectures with nonlinear transitions are well-suited for solving the group word problem, because in theory and practice, such architectures can learn the group word problem to high accuracy with a single layer. Other literature has explored the value of matrix-valued states (Beck et al., 2024; Grazzi et al., 2025). However, in Proposition˜4 below, we show that neither nonlinearity nor matrix-valued states are needed to understand or solve the group word problem. Instead, the problem can be formulated as an LDS with vector-valued states and input-dependent transition matrices.
@@
By Cayley’s theorem, any finite group $G$ can be embedded in a symmetric group $S_{D}$, for some $D\leq|G|$. Therefore, by choosing the initial state $x_{0}\in\mathbb{R}^{D}$ to have $D$ distinct entries (a “vocabulary” of size $D$), we can use the tabular representation of permutations (Artin, 2011, eq. 1.5.2) to represent an element of $S_{D}$ as $x_{t}$ (by a permutation of the elements of $x_{0}$). We can also choose $A_{t+1}\in\mathbb{R}^{D\times D}$ to be the permutation matrix corresponding to the embedding of $g_{t+1}$ in $S_{D}$, since any element of $S_{D}$ can be represented as a $D\times D$ permutation matrix (e.g., see Figure˜2B). Consequently $x_{t}=A_{t}A_{t-1}\ldots A_{2}A_{1}x_{0}$ is an embedding of an element of $G$ in $S_{D}$ in the tabular representation. In fact, $x_{t}\in\mathbb{R}^{D}$ represents the running product $g_{1}g_{2}\ldots g_{t-1}g_{t}$, which is precisely the goal of the group word problem. ∎
@@
Figure 2: A single Newton iteration solves the $S_{5}$ group word problem, whereas the number of iterations required for the other methods increases with sequence length. We consider the task of evaluating the product of $S_{5}$ group elements. A: The group word problem can be expressed as an LDS with input-dependent state-transition matrices. B: An example input-dependent transition matrix $A_{t}$ for permutation (1 5 2 4 3)(1\ 5\ 2\ 4\ 3), in cycle notation. C: For each fixed-point method and a range of sequence lengths, $T$, we compute the median (over ten random seeds) number of fixed-point iterations to converge (top) and the median wall-clock time (bottom). While a single Newton iteration is sufficient to solve the $S_{5}$ problem, the number of iterations required for the other methods increases with the sequence length.
@@
Though we have cast the group word problem as a time-varying LDS with $f_{t+1}(x_{t})=A_{t+1}x_{t}$, we can still evaluate this recursion with any of the fixed-point methods described above. Since the dynamics are linear, the Newton iteration corresponds to evaluating the LDS with a parallel scan, and it converges in one iteration. While other methods would require more iterations to converge, they could still be more efficient in wall-clock time, since they use less memory and compute per iteration. However, we can use the Jacobian approximation error of the different fixed-point methods (Lemma˜1) to get a sense if the other fixed-point methods are likely to excel in this setting.
The state transition matrices of the group word problem are permutation matrices with spectral norm one, and so $\mathrm{diff}(\mathcal{A}_{J})=1$. Furthermore, since with high probability there will be a state transition matrix with no non-zero entries on the diagonal, it follows that $\mathrm{diff}(\mathcal{A}_{QN})=1$ while $\mathrm{diff}(\mathcal{A}_{P})=2$. Since we would expect to need $\mathrm{diff}(\mathcal{A})<1$ for a fixed-point method $\mathcal{A}$ to be effective, our theoretical analysis in Section˜3 suggests that none of the fixed-point methods other than Newton will be effective on the group word problem.
@@
We test this hypothesis with a simple experiment simulating the $S_{5}$ word problem, a standard problem in the sequence modeling literature (Merrill et al., 2024; Grazzi et al., 2025). In this setting, Figure˜2 shows that quasi-Newton, Picard, and Jacobi iterations require nearly $T$ iterations to converge. On the other hand, we see that Newton’s method solves the $S_{5}$ word problem with just one fixed-point iteration, as expected since the true dynamics are linear. The speed-up is also apparent in the wall-clock time comparison, where we see that Newton is faster than other methods, regardless of $T$.
4.2 Case Study #2: Picard iterations struggle to parallelize RNNs